# ETH-monotonicity and the black hole singularity

### 11 Mar 2026 v3

@@ Appendix B Further numerical results
Figure 6(a) is plots of the function $f(\bar{E})$ for $\Delta=\Delta_{-}=3/2,l=0$, corresponding to $m^{2}=-15/4$. We have used (2.13) for this calculation. We again find that $f(\bar{E})$ is a monotonic function of $\bar{E}$, in general. In this case, the quasiparticles will appear at $\omega=\Delta_{-}+2n,n\in\mathbb{N}_{0}$. Figure 6(b) is plots of the $f$-function with varying $\omega$ for $\Delta_{-}=3/2,l=0$ with different fixed values of $r_{h}$. We can see the quasiparticle peak prominently for smaller value of $r_{h}$. We can also see the exponential fall-off of $e^{-\beta\omega/4}$.
@@
Figure 7 are numerical results of the 3-d CFT from $(3+1)$-D asymptotically AdS black hole. We again find that the scalar operator satisfies ETH-monotonicity.
@@
Figure 8: Plots of $\bar{E}f^{\prime}(\bar{E})/f(\bar{E})=\mu f^{\prime}(\mu)/f(\mu)$ as $\bar{E}\to 0$ considering different values of $l$ for (a) 4d4-d holographic CFT with $\Delta=\Delta_{-}=3/2$ and (b) 3d3-d holographic CFT with $\Delta=\Delta_{+}=5/4$. The limiting value $\kappa$ is again given by $\kappa=\kappa_{0}+d_{\kappa}l$ with $\kappa_{0}=0.5$.

### 25 Mar 2026 v4

@@ Appendix B Further numerical results
Figure 6(a) is plots of the function $f(\bar{E})$ for $\Delta=\Delta_{-}=3/2,l=0$, corresponding to $m^{2}=-15/4$. We have used (2.13) for this calculation. We again find that $f(\bar{E})$ is a monotonically increasing function of $\bar{E}$, in general. In this case, the quasiparticles will appear at $\omega=\Delta_{-}+l+2n,n\in\mathbb{N}_{0}$. Figure 6(b) is plots of the $f$-function with varying $\omega$ for $\Delta_{-}=3/2,l=0$ with different fixed values of $r_{h}$. We can see the quasiparticle peak prominently for smaller value of $r_{h}$. We can also see the exponential fall-off of $e^{-\beta\omega/4}$.
@@
Figure 7 are numerical results for the 3-d holographic CFT from $(3+1)$-D asymptotically AdS black hole. We again find that the scalar operator satisfies ETH-monotonicity.
@@
Figure 8(a) is plots of $\bar{E}f^{\prime}(\bar{E})/f(\bar{E})$ for $\Delta=\Delta_{-}=3/2$ with different values of $l$ for 4-d holographic CFT. Figure 8(b) is plots of $\bar{E}f^{\prime}(\bar{E})/f(\bar{E})$ for $\Delta=\Delta_{-}=5/4$ with different values of $l$ for 3-d holographic CFT. The wild variation for $l=0,\;\omega=1.2$ is because of the quasiparticle peak at $\omega=5/4=1.25$.
Figure 8: Plots of $\bar{E}f^{\prime}(\bar{E})/f(\bar{E})=\mu f^{\prime}(\mu)/f(\mu)$ as $\bar{E}\to 0$ considering different values of $l$ for (a) 4d4-d holographic CFT with $\Delta=\Delta_{-}=3/2$ and (b) 3d3-d holographic CFT with $\Delta=\Delta_{-}=5/4$. The limiting value $\kappa$ is again given by $\kappa=\kappa_{0}+d_{\kappa}l$ with $\kappa_{0}=0.5$.